# preprint: IFT-UAM/CSIC-25-123

### v1

@@ Abstract
Introduction. Quantum complexity has become a central concept at the intersection of quantum information, condensed-matter theory, and high-energy physics. One of the most influential approaches is Nielsen complexity [Nielsen:2005mkt, Nielsen:2006cea, Dowling:2006tnk], which reformulates circuit complexity as a geometric problem. In this framework, a circuit corresponds to a curve on the unitary group manifold, and the complexity is given by the length of the minimal geodesic connecting the chosen unitary operator with the identity operator [Jefferson:2017sdb, Khan:2018rzm, Hackl:2018ptj, Chapman:2018hou, Bhattacharyya:2018bbv, Guo:2018kzl, Bernamonti:2019zyy, Bernamonti:2020bcf, Caceres:2019pgf, Ruan:2020vze]. A complementary measure, Krylov complexity [Parker:2018yvk], quantifies operator growth under Heisenberg evolution. Starting from a seed operator, one constructs an orthonormal Krylov basis via the Lanczos algorithm, thereby mapping the dynamics to a one-dimensional hopping problem along the Krylov chain. Krylov complexity directly measures the spread of the initial operator in Krylov space and provides a quantitative probe of operator growth, scrambling, and chaos [Barbon:2019wsy, Avdoshkin:2019trj, Rabinovici:2020ryf, Jian:2020qpp, Dymarsky:2021bjq, Hornedal:2022pkc, Balasubramanian:2022tpr, Erdmenger:2023wjg, Caputa:2024vrn, Baggioli:2024wbz, Craps:2024suj, Huh:2024ytz]. For comprehensive reviews, see [Nandy:2024htc, Baiguera:2025dkc, Rabinovici:2025otw].
Despite their different origins, Nielsen complexity and Krylov complexity both have deep ties to holography. Nielsen’s geometric framework has motivated holographic proposals linking circuit complexity to geometric quantities that capture the growth of the black hole interior [Susskind:2014rva, Stanford:2014jda, Brown:2015bva, Brown:2015lvg, Cai:2016xho, Belin:2021bga, Belin:2022xmt, Carrasco:2023fcj, Jorstad:2023kmq, Jiang:2023jti, Caceres:2023ziv, Myers:2024vve, Arean:2024pzo, Jiang:2025qai, Miyaji:2025yvm, Miyaji:2025jxy, Caceres:2025myu]. Meanwhile, in models of two-dimensional gravity, Krylov state complexity provides a concrete realization of the proposed correspondence between complexity and size of black hole interiors [Lin:2022rbf, Rabinovici:2023yex, Heller:2024ldz, Balasubramanian:2024lqk, Heller:2025ddj, Ambrosini:2025hvo, Fu:2025kkh]. Although the possible relation between Nielsen complexity and Krylov complexity has attracted significant recent interest [Lv:2023jbv, Caputa:2021sib, Craps:2023ivc, Aguilar-Gutierrez:2023nyk], a direct and general correspondence remains elusive. These developments raise a central question: are Nielsen complexity and Krylov complexity fundamentally related? The aim of this work is to provide such an explicit correspondence.
@@
with the lower bound saturated at $\alpha=0$. In part B of the Supplemental Material, we analytically verify that this bound also holds for the Berger sphere with arbitrary penalties. We thus find that, in all models that we investigated, the straight-line trajectory (7) remains a locally minimal geodesic for any precursor operator with $z\leq\tfrac{\pi}{2}$. This motivates us to expect that, also for the SYK model with penalty schedules of interest, the straight-line trajectory (7) will be a locally minimal geodesic for a finite range of $z$

### v2

@@ Abstract
Introduction. Quantum complexity has become a central concept at the intersection of quantum information, condensed-matter theory, and high-energy physics. One of the most influential approaches is Nielsen complexity [Nielsen:2005mkt, Nielsen:2006cea, Dowling:2006tnk], which reformulates circuit complexity as a geometric problem. In this framework, a circuit corresponds to a curve on the unitary group manifold, and the complexity is given by the length of the minimal geodesic connecting the chosen unitary operator with the identity operator [Jefferson:2017sdb, Khan:2018rzm, Hackl:2018ptj, Chapman:2018hou, Bhattacharyya:2018bbv, Guo:2018kzl, Bernamonti:2019zyy, Bernamonti:2020bcf, Caceres:2019pgf, Ruan:2020vze, Haque:2024ldr]. A complementary measure, Krylov complexity [Parker:2018yvk], quantifies operator growth under Heisenberg evolution. Starting from a seed operator, one constructs an orthonormal Krylov basis via the Lanczos algorithm, thereby mapping the dynamics to a one-dimensional hopping problem along the Krylov chain. Krylov complexity directly measures the spread of the initial operator in Krylov space and provides a quantitative probe of operator growth, scrambling, and chaos [Barbon:2019wsy, Avdoshkin:2019trj, Rabinovici:2020ryf, Jian:2020qpp, Dymarsky:2021bjq, Hornedal:2022pkc, Balasubramanian:2022tpr, Erdmenger:2023wjg, Caputa:2024vrn, Baggioli:2024wbz, Craps:2024suj, Huh:2024ytz]. For comprehensive reviews, see [Nandy:2024htc, Baiguera:2025dkc, Rabinovici:2025otw].
Despite their different origins, Nielsen complexity and Krylov complexity both have deep ties to holography. Nielsen’s geometric framework has motivated holographic proposals linking circuit complexity to geometric quantities that capture the growth of the black hole interior [Susskind:2014rva, Stanford:2014jda, Brown:2015bva, Brown:2015lvg, Cai:2016xho, Belin:2021bga, Belin:2022xmt, Carrasco:2023fcj, Jorstad:2023kmq, Jiang:2023jti, Caceres:2023ziv, Myers:2024vve, Arean:2024pzo, Jiang:2025qai, Miyaji:2025yvm, Miyaji:2025jxy, Caceres:2025myu]. Meanwhile, in models of two-dimensional gravity, Krylov state complexity provides a concrete realization of the proposed correspondence between complexity and size of black hole interiors [Lin:2022rbf, Rabinovici:2023yex, Heller:2024ldz, Balasubramanian:2024lqk, Aguilar-Gutierrez:2024nau, Xu:2024gfm, Heller:2025ddj, Ambrosini:2025hvo, Fu:2025kkh]. Although the possible relation between Nielsen complexity and Krylov complexity has attracted significant recent interest [Lv:2023jbv, Caputa:2021sib, Craps:2023ivc, Aguilar-Gutierrez:2023nyk], a direct and general correspondence remains elusive. These developments raise a central question: are Nielsen complexity and Krylov complexity fundamentally related? The aim of this work is to provide such an explicit correspondence.
@@
with the lower bound saturated at $\alpha=0$. In part B of the Supplemental Material, we analytically verify that this bound also holds for the Berger sphere with arbitrary penalties. We thus find that, in all models that we investigated, the straight-line trajectory (7) remains a locally minimal geodesic for any precursor operator with $z\leq\tfrac{\pi}{2}$. This motivates us to expect that, also for the SYK model with penalty schedules of interest, the straight-line trajectory (7) will be a locally minimal geodesic for a finite range of $z$.